OpenStudy (anonymous):

convert to polar y=x^2

OpenStudy (anonymous):

\[x^2+y^2=r^2~~\iff~~x^2=r^2-y^2~~\iff~~x^2=r^2-r^2\sin^2\theta\] So \[r\sin\theta=r^2\left(1-\sin\theta\right)\]

OpenStudy (anonymous):

thanks

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."